# LLM-Assisted Request Mutation for Whitebox REST API Testing

### v2

@@ 3.1. High-Level Walkthrough
1:$program$ to test, search time $T$, $model$ to query for test cases, and minimum count of times to query per iteration $M$.
2:a set of test cases exercising the program that maximizes the coverage.
3:$apis\leftarrow\text{{GetApis}}(program)$
4:$testCases\leftarrow\text{{RandomTestCases}}(apis)$
5:$archive\leftarrow\text{{UpdateCoverage}}(\emptyset\cup testCases,covPts)$
6:while $timeElapsed<T$ do
7:    $chosenTarget\leftarrow\text{{ChooseTargetByUncover}}(archive)$
8:    $chosenTestCase\leftarrow\text{{ChooseTestCases}}(testCases,chosenTarget,archive)$
9:    $relatedCode\leftarrow null$
[…]
11:       $relatedCode\leftarrow\text{{GetRelatedCode}}(chosenTarget)$
12:    end if
[…]
14:    $llmTimes\leftarrow 0$
15:    if $relatedCode\neq null$ then
16:       $llmTimes\leftarrow\text{{Max}}(\frac{upToNTimes}{2},M)$
17:    end if
18:    $totalTimes\leftarrow\text{{Max}}(upToNTimes,llmTimes)$
19:    for $i\leftarrow 1$ to $totalTimes$ do
20:       if $i\leq llmTimes$ then
21:          $newTestCase\leftarrow\text{{LLMAssitedMutate}}(chosenTestCase,apis,\newline relatedCode,model)$
22:       else
23:          $newTestCase\leftarrow\text{{Mutate}}(chosenTestCase,apis)$
24:       end if
25:       $evaluateResult\leftarrow\text{{calculateCoverage}}(archive,chosenTestCase,\newline newTestCase)$
26:       if $\text{{CoverNotFewer}}(evaluateResult)$ then
27:          $archive\leftarrow\text{{UpdateCoverage}}(archive\cup newTestCase)$
28:          $chosenTestCase\leftarrow\text{{UpdateCurrentTestCases}}(newTestCase)$
29:          $testCases\leftarrow\text{{UpdateTestCases}}(testCases\cup newTestCase)$
30:       end if
31:    end for
32:end while
33:return $archive$
@@
Algorithm 1 shows MioHint’s high-level algorithm. MioHint is an extension of the MIO (Many Independent Objectives) algorithm. It incorporates a large language model (LLM) to improve the mutation process for generating API test cases, with the aim of maximizing code coverage. Parts of the algorithm that remain unchanged from MIO are indicated in grey.
@@
The algorithm begins by extracting the set of APIs from the program under test using the function GetApis(program) (Line 1). Then, it samples random test cases according to the API definition (Line 2), calculates its coverage, and updates the coverage to the archive (Line 3) as in Algorithm MIO. After the target is chosen, MioHint extracts code related to the chosen target when the target is of type branch or method replacement (Line 8-9). We will further elaborate on how we perform our related code extraction in the next section as a vital part of our algorithm. If relatedCode exists, llmTimes is set to the maximum of half the number of mutations and a minimum count $M$ (Lines 14-15). We set M to 2 in our experiment because of the randomness of response from the LLM. We also adapt the total mutation times to LLM-assisted mutation times to ensure the LLM-assisted mutation performs at least twice (Lines 16). We replace half of the original mutation times of vanilla mutation with our proposed LLM-assisted mutation (Lines 19). The subsequent evaluation steps of the algorithm align with Algorithm MIO. If the new test case covers not fewer targets than the current one, update the current test case (chosenTestCase), as well as the archive and testCases (Lines 23-28).
3.2. LLM-assited Mutation
@@
Definition of Hard-to-cover Targets. We define hard-to-cover targets as those selected during the search process that fall within the scope of our analysis. These targets are characterized by the inefficiency of naive random mutation in generating the correct mutants needed to cover them, i.e., numerous iterations of mutation often fail to produce a successful mutant.
@@ 3.2.1. Related Code Extraction
Figure 6 provides an example of the code we extract. Initially, we process the specific file containing the target line boolean matches = matcher.matches(); and the encompassing function public boolean isHashValid(String toValidate) {...} to form the local context. Notably, for function code, we retain the annotations of function to leverage the LLM’sunderstanding of natural language. For global context, we perform cross-file analysis within the entire repository to get the value expansion including def-use analysis and called function expansion. The def-use analysis consisted of two parts. Part 1 of the def-use analysis is conducted within the functions at the call chain of the focal function, aiming to identify the value of the passing argument toValidate. When calling function isHashValid(key), parameter key is passed, thus toValidate = key, and the definition of the parameter key is key = request.getKey(). In Part 2 of the def-use analysis, the process iteratively seeks the definition of variables used in the target. First, find the definition of the function variable matcher which is used at the focal line. Then find the definition of the class variable PATTERN imported by the definition of matcher. Finally, expand the definition of the class variable GUID_HASH_PATTERN imported by the definition of PATTERN. This iterative approach allows for a comprehensive understanding of the data flow and dependencies associated with the variables involved.
@@
$targetStatement$
a chain of definition statements $defUseChain$
@@
function analyzeVariable($variable$)
@@
    if $defStmt$ not in $defUseChain$ then
@@
        for each $nv$ in $newVariables$ do
           analyzeVariable($nv$)
@@
function analyzeCallers($function$)
@@
return $defUseChain$
We present details of def-use analysis in Algorithm 2. This algorithm constructs a chain of definition statements of used variables for a given target statement according to the def-use relationship. The main procedure initializes an empty set $defUseChain$ and extracts variables from the $targetStatement$. For each variable, it calls the $analyzeVariable$ function, which recursively finds the definition statements of the variables and adds them to the $defUseChain$. The main procedure also identifies the enclosing function of the $targetStatement$ and calls the $analyzeCallers$ function with it. The $analyzeCallers$ function identifies all callers of the given function, extracts parameters of the call statement, and recursively analyzes these parameters to extend the $defUseChain$. The $analyzeCallers$ function analyzes the caller’s caller iteratively until includes all functions in the call chain. The algorithm ensures that all relevant definition statements are included in the $defUseChain$ by the end of its execution.
Analysis Scope. For each hard-to-cover target, we analyze only a small portion of the files within the entire repository, which helps to alleviate the analysis overhead. In the local context, the analysis scope is limited to the files that contain the target. In the global context, the analysis scope extends to the files along the call chain leading to the surrounding function of the target. We preserve the paths of the files containing the caller functions in the call chain and switch the file under analysis to the one containing the caller function when analyzing call parameters.
@@ 3.2.2. Prompt Construction
Having retrieved the target along with related code and REST call actions to be mutated, we construct a prompt following a structured approach. As illustrated in Figure 7, we begin with a task-specific instruction, labeled (1). The mutation task is then decomposed into three sub-tasks to enhance the reasoning process through Chain-of-Thought (CoT). This decomposition involves selecting the most relevant REST call action, locating the specific type of parameter along with its field, and finally, revising the selected field, labeled (2).
To further refine the mutation process, we increase attention to dependent conditions within the control flow, labeled (3). Subsequently, we populate the template with previously extracted information, including the target, line of code, def-use chain, called function definition, function code, and REST call action, labeled (4). We specify the desired output format by providing illustrative examples, labeled (5). Finally, we incorporate hints generated from previous runs along with their execution feedback to guide the mutation process, leveraging in-context learning to learn from historical feedback, labeled (6).
@@ 3.2.3. Test Case Generation
We query the LLM using the prompt constructed in the previous section and post-process its response to extract a hint represented as a JSON string. This hint includes information on which REST call action to revise, the type of parameter to modify, the specific field to update, and the new value for that field. After applying the hint to the original test case, we get a mutated test case that aims to cover the selected target as the output of LLM-assisted mutation. The new test case will be evaluated soon and updated to the population according to its evaluation result. If the new test case is evaluated not worse, later mutation will be applied to it. The evaluation result will be preserved as feedback to the next LLM-assisted mutation for the same target.
@@ 4. Evaluation
  • RQ1: What’s the performance on line coverage, target coverage, and mutation hit rate of our method?
  • RQ2: What’s the contribution of value expansion in our method to the overall performance?
  • RQ3: What’s the runtime overhead introduced by our method?
@@ 4.1. Evaluation Setup
  • Average Execution Time per Test, this metric represents the execution time of every run of test case mutation and evaluation.
Evaluation Datasets. We use real-world programs from the EMB (EMB, [n. d.]) corpus as the evaluation benchmarks. EMB is a well-maintained open-source corpus of Web APIs (including REST, GraphQL, and RPC APIs). The detail of the real-world benchmark programs we used is shown in Table 1, including the number of source files, lines of code, and the number of REST endpoints in each API. These statistics only account for the business logic code and exclude third-party libraries (e.g., HTTP servers).
@@
Some SUTs were excluded due to objective factors. Specifically, the features-service encountered a runtime error. Additionally, all targets within the cwa-verification, scout-api, and session-service are related to database operations or configurations, which modify requests can not address. Furthermore, the ocvn, proxyprint, market, and reservations-api do not contain any targets within the scope of analysis during the search process.
Experiment Settings. Each fuzzing session runs for 1 hour. To account for the randomness of search-based fuzzing, each experiment was repeated 10 times. With 3 settings and 16 SUTs, required 3 × 16 × 10 = 480 hours, i.e., 20 days of computation. We conducted our evaluations on the machine equipped with Intel Xeon Gold 5218R CPU with 10 cores, using Ubuntu 20.04.6 LTS as the operating system. For LLM query, We use public APIs provided by OpenAI with GPT-4o (OpenAI, 2024).
@@
4.2. RQ1: Line Coverage, Target Hit Rate, and Mutation Hit Rate of baseline and our method
Results are compared in term of line coverage, target coverage, and mutation hit rate. Metric line coverage evaluates the overall performance impact of our method and baseline. Metric target coverage presents our method’s efficiency in covering hard-to-cover targets. Metric mutation hit rate evaluates the accuracy of our LLM-assisted mutation compared with naive mutation. Table 2 shows the results in detail for our method compared to the baseline. We follow the statistical guidelines from  (Arcuri and Briand, 2011), reporting $p$-values of Mann-Whitney-Wilcoxon U tests and Vargha-Delaney standarized $\hat{A}_{12}$ effect sizes.
For metric line coverage, our method improves line coverage for most SUTs. Results are statistical significant for 4 SUTs, with no statistically worse results. On these APIs improvement are either ”medium” (e.g., +2.9% for catwatch and +4.4% for genome-nexus) or ”large” (e.g., +11.1% for languagetool and +18.0% for rest-scs). Overall, our method achieved an average increase in line coverage of 4.95% (from 48.51% to 53.46%). This indicates that our method contributes not only to the selected target but also to the overall performance of the search algorithm.
@@ 4.2. RQ1: Line Coverage, Target Hit Rate, and Mutation Hit Rate of baseline and our method
To elucidate the relationship between line coverage and target coverage, and to explain anomalies such as service pay-publicapi where target coverage increases from 0% to 100% without a corresponding change in line coverage, it is important to consider the underlying factors. Line coverage is strongly correlated with the total number of covered targets, and what we do is design a precise mutation strategy to increase the coverage of selected targets. Thus, when the total number of selected targets is low, even if the target coverage is high, the overall line coverage may not exhibit a significant increase because the total number of covered targets is low. For example, pay-publicapi only has 2 selected targets. This phenomenon can be attributed to the nature of the target selection strategy, which is governed by the underlying search-based algorithm, MIO. Consequently, there will be a situation where the target coverage will increase a lot but the line coverage will not increase much. Similarly, the reason why the service gestaohospital achieves almost equal line coverage under both settings can be attributed to the low number of targets, which is 5 or 6.
Mutation hit rate is a crucial metric for assessing the quality of test cases generated by mutation. Our method also shows significant improvements in this metric. For example, the mutation hit rate for genome-nexus increased from 0.18% to 12.23%, and for rest-scs, it rose from 0.42% to 28.16%. Overall, our method achieved an average increase in mutation hit rate of 21.82 percentage points (from 0.35% to 22.17%). This indicates that the accuracy of our LLM-assisted mutation is 67$\times$ higher than that of the vanilla mutation.
@@
Answer to RQ1: The experimental data shows that our method consistently outperforms the baseline across several metrics: average line coverage (53.46% vs. 48.51%), an increase of 4.95%; target coverage (57.54% vs. 9.82%), more than half targets; and mutation hit rate (67$\times$).
@@
4.3. RQ2: The effectiveness of value expansion
To demonstrate the effectiveness of our value expansion technique, we conduct an ablation study at value expansion and evaluate it using the same metrics in RQ1. Table 3 shows the results in detail for the performance between three different configurations: the baseline, our proposed method, and our method without variable expansion (no VE).
After disabling value expansion, line coverage decreases by 1.5%, target coverage decreases by 8.5%, and the mutation hit rate drops from 22.2% to 12.9%, which is approximately half of the original rate. The decrease in performance across all metrics highlights the effectiveness of our value expansion. This indicates that with our value expansion, an additional 8.5% of the hard-to-cover targets can be addressed; these targets are more challenging than the 48.5% targets covered without value expansion as they need more contextual information to build helpful mutation hints.
Answer to RQ2: Our value expansion brings 1.5% of line coverage improvement, and enables coverage of 8.5% additional challenging targets that require contextual mutation hints.
4.4. RQ3: Runtime Overhead

### v3

@@ 3.1. High-Level Walkthrough
Algorithm 1 shows MioHint’s high-level algorithm. MioHint is an extension of the MIO (Many Independent Objectives) algorithm. It incorporates an LLM to improve the mutation process for generating API test cases, with the aim of maximizing code coverage. For clarity, the parts of the algorithm that remain unchanged from the original MIO are shown in grey.
We designed MioHint by preserving the core framework of the MIO algorithm, including its target selection, test case evaluation, and archiving mechanisms, and integrating our LLM-assisted mutation module. This synergistic design is motivated by the crucial balance between efficiency and effectiveness. The fast random mutation is highly efficient for exploring the general test space and covering readily reachable targets. The more resource-intensive LLM is then strategically deployed as a specialist, invoked to overcome hard-to-cover targets selected by MIO that cause the search process to stagnate. The test case generated by the LLM is then reintegrated into the main search loop: it is evaluated by the MIO fitness metric coverage, and if deemed high-quality, it is archived to serve as a new, promising seed for subsequent random mutations, which further expand the exploration. This principled strategy maximizes overall testing performance by combining the broad, efficient exploration of the search algorithm with the deep, targeted problem-solving of the LLM.
@@
The algorithm begins by extracting the set of APIs from the program under test using the function GetApis(program) (Line 1). Then, it samples random test cases according to the API definition (Line 2), calculates its coverage, and updates the coverage to the archive (Line 3) as in Algorithm MIO. After the target is chosen, MioHint extracts code related to the chosen target when the target is of type branch or method replacement (Line 8-9). We will further elaborate on how we perform our related code extraction in the next section as a vital part of our algorithm. If relatedCode exists, llmTimes is set to the maximum of half the number of mutations and a minimum count $M$ (Lines 14-15). We set M to 2 in our experiment because of the randomness of responses from the LLM. We also adapt the total mutation times to LLM-assisted mutation times to ensure the LLM-assisted mutation performs at least twice (Lines 16). We replace half of the original mutation times of vanilla mutation with our proposed LLM-assisted mutation (Lines 19). The subsequent evaluation steps of the algorithm align with Algorithm MIO. If the new test case covers not fewer targets than the current one, update the current test case (chosenTestCase), as well as the archive and testCases (Lines 23-28).
3.2. LLM-assisted Mutation
@@
Definition of Hard-to-cover Targets. We define hard-to-cover targets as those selected during the MIO algorithm search process that fall within the scope of our analysis. Note that we did not perform additional filtering for these targets. These targets are characterized by the inefficiency of naive random mutation in generating the correct mutants needed to cover them, i.e., numerous iterations of mutation often fail to produce a successful mutant.
@@ 3.2.1. Related Code Extraction
Figure 6 provides an example of the code we extract. Initially, we process the specific file containing the target line boolean matches = matcher.matches(); and the encompassing function public boolean isHashValid(String toValidate) {...} to form the local context. Notably, for function code, we retain the annotations of the function to leverage the LLM’s understanding of natural language. For global context, we perform cross-file analysis within the entire repository to get the value expansion including def-use analysis and called function expansion. The def-use analysis consisted of two parts. Part 1 of the def-use analysis is conducted within the functions at the call chain of the focal function, aiming to identify the value of the passing argument toValidate. When calling function isHashValid(key), parameter key is passed, thus toValidate = key, and the definition of the parameter key is key = request.getKey(). In Part 2 of the def-use analysis, the process iteratively seeks the definition of variables used in the target. First, find the definition of the function variable matcher which is used at the focal line. Then find the definition of the class variable PATTERN imported by the definition of matcher. Finally, expand the definition of the class variable GUID_HASH_PATTERN imported by the definition of PATTERN. This iterative approach allows for a comprehensive understanding of the data flow and dependencies associated with the variables involved.
@@
0:  targetStatementtargetStatement
0:  a chain of definition statements defUseChaindefUseChain
@@
  function analyzeVariable(variablevariable)
@@
    if defStmtdefStmt not in defUseChaindefUseChain then
@@
      for each nvnv in newVariablesnewVariables do
        analyzeVariable(nvnv)
@@
  function analyzeCallers(functionfunction)
@@
  return defUseChaindefUseChain
We present details of def-use analysis in Algorithm 2. This algorithm constructs a chain of definition statements of used variables for a given target statement according to the def-use relationship. The main procedure initializes an empty set $defUseChain$ and extracts variables from the $targetStatement$. For each variable, it calls the $analyzeVariable$ function, which recursively finds the definition statements of the variables and adds them to the $defUseChain$. The main procedure also identifies the enclosing function of the $targetStatement$ and calls the $analyzeCallers$ function with it. The $analyzeCallers$ function identifies all callers of the given function, extracts parameters of the call statement, and recursively analyzes these parameters to extend the $defUseChain$. The $analyzeCallers$ function analyzes the caller’s caller iteratively until it includes all functions in the call chain. The algorithm ensures that all relevant definition statements are included in the $defUseChain$ by the end of its execution.
Analysis Scope. The analysis is narrowly focused, helping minimize overhead. For each hard-to-cover target, the analysis scope of local context is restricted to the file containing the target. The global context extends this to include files traced along the call chain of the target’s enclosing function.
@@ 3.2.2. Prompt Construction
Having retrieved the target along with related code and REST call actions to be mutated, we construct a prompt following a structured approach. There are six input features in the prompt, each providing orthogonal information. As illustrated in Figure 7, we begin with a task-specific instruction, labeled (1). The first feature defines the task, enabling LLM to understand what is required to do. The second feature adopts a chain-of-thought (CoT) approach to stepwise locate the object to be mutated, from the request down to the specific parameter. Specifically, this CoT involves selecting the most relevant REST call action, locating the specific type of parameter along with its field, and finally, revising the selected field, labeled (2). To further refine the mutation process, the third feature increases attention to control flow conditions to address implicit flows, labeled (3). Subsequently, we populate the template with previously extracted information, including the target, line of code, def-use chain, called function definition, function code, and REST call action, labeled (4). Specifically, in the fourth feature, the Target is specified as a string (as shown in Figure 5), indicating the target’s type, class, line, position, and expected value. LineCode is used to locate the target within the FunctionCode; the DefUseChain and CalledFunctionDefinition provide data flow relationships between the target and request; and RestCallActions represent the detailed content of the requests that can be selected for mutation. The fifth feature defines the output format as well-defined JSON by providing illustrative examples for consistent output parsing, labeled (5). Finally, the sixth feature incorporates hints generated from previous runs along with their execution feedback to guide the mutation process. Specifically, it leverages in-context learning to learn from historical feedback, thereby avoiding repeatedly generating incorrect mutations, labeled (6). The absence of any of these steps can introduce errors in LLM’s reasoning process.
@@ 3.2.3. Test Case Generation
We query the LLM using the prompt constructed in the previous section. Its response is then post-processed to extract a mutation hint, which is represented as a JSON string. This hint specifies the REST call action to revise, the parameter type to modify, the specific field to update, and the new value for that field. Subsequently, the hint is applied to the original test case to generate a mutated version. The resulting mutated test case, designed to cover the selected target, is then evaluated. If the evaluation shows no coverage degradation, it is added to the population and becomes eligible for further mutation. Furthermore, the evaluation result is preserved and used as feedback for subsequent LLM-assisted mutation attempts on the same target.
@@ 4. Evaluation
  • RQ1: How effective is MioHint in generating test cases to cover the SUT?
  • RQ2: To what extent does the value expansion component improve the test generation effectiveness of MioHint?
  • RQ3: What is the runtime efficiency of MioHint?
@@ 4.1. Evaluation Setup
  • Average execution time per test, this metric represents the execution time of every run of the test case mutation and evaluation.
Evaluation Datasets. We use real-world programs from the EMB (EMB, ) corpus as the evaluation benchmarks. EMB is a well-maintained open-source corpus of Web APIs (including REST, GraphQL, and RPC APIs). The detail of the real-world benchmark programs we used is shown in Table 1, including the number of source files, lines of code, and the number of REST endpoints in each API. These statistics only account for the business logic code and exclude third-party libraries (e.g., HTTP servers).
@@
Some SUTs were excluded due to objective factors. Specifically, the features-service encountered a runtime error. Additionally, all targets within the cwa-verification, scout-api, and session-service are related to database operations or configurations, which modify requests that can not be addressed. Furthermore, the ocvn, proxyprint, market, and reservations-api do not contain any targets within the scope of analysis during the search process.
Experiment Settings. Each fuzzing session runs for 1 hour. To account for the randomness of search-based fuzzing, each experiment was repeated 10 times. With 3 settings and 16 SUTs, it required 3 × 16 × 10 = 480 hours, i.e., 20 days of computation. We conducted our evaluations on the machine equipped with an Intel Xeon Gold 5218R CPU with 10 cores, using Ubuntu 20.04.6 LTS as the operating system. For LLM query, we use public APIs provided by OpenAI with GPT-4o (OpenAI, 2024).
@@
4.2. RQ1: Effectiveness on Covering SUT
To assess the effectiveness of our LLM-assisted design, we conduct a comparative evaluation against a baseline across three key metrics: line coverage, target coverage, and mutation hit rate. Specifically, line coverage serves as a general indicator of the overall code coverage achieved by both our method and the baseline. Higher coverage values indicate that MioHint is more effective at generating test cases that thoroughly exercise the SUTs. Furthermore, target coverage specifically quantifies our method’s effectiveness in covering hard-to-cover targets. Lastly, mutation hit rate assesses the accuracy and effectiveness of our LLM-assisted mutation strategy compared to naive mutation. Table 2 shows the results in detail for our method compared to the baseline. We follow the statistical guidelines from  (Arcuri and Briand, 2011), reporting $p$-values of Mann-Whitney-Wilcoxon U tests and Vargha-Delaney standarized $\hat{A}_{12}$ effect sizes.
For metric line coverage, our method improves line coverage for most SUTs. Results are statistically significant for 4 SUTs, with no statistically worse results. On these APIs improvement are either ”medium” (e.g., +2.9% for catwatch and +4.4% for genome-nexus) or ”large” (e.g., +11.1% for languagetool and +18.0% for rest-scs). Overall, our method achieved an average increase in line coverage of 4.95% (from 48.51% to 53.46%). This indicates that our method contributes not only to the selected target but also to the overall performance of the search algorithm.
@@ 4.2. RQ1: Effectiveness on Covering SUT
To clarify the relationship between line coverage and target coverage, and to explain anomalies such as the case of the service pay-publicapi, where target coverage increases from 0% to 100% without a corresponding rise in line coverage, it is important to consider the underlying factors. Line coverage is closely related to the total number of covered targets. Our mutation strategy is designed to maximize the coverage of selected targets, but when the number of targets is low, even full target coverage may not significantly increase overall line coverage. For example, pay-publicapi has only two selected targets. This is due to the target selection strategy determined by the underlying MIO search-based algorithm. As a result, target coverage can increase substantially while line coverage remains largely unchanged. Similarly, for the service gestaohospital, the nearly identical line coverage under both settings is also due to the small number of targets, which is five or six.
[…]
@@
[…]
@@
4.3. RQ2: Effectiveness of Value Expansion
To demonstrate the effectiveness of our value expansion technique, we conduct an ablation study on value expansion and evaluate it using the same metrics in RQ1. Table 3 shows the results in detail for the performance between three different configurations: the baseline, our proposed method, and our method without variable expansion (no VE).
After disabling value expansion, line coverage decreases by 1.5%, target coverage decreases by 8.5%, and the mutation hit rate drops from 22.2% to 12.9%, which is approximately half of the original rate. The decrease in performance across all metrics highlights the effectiveness of our value expansion. This indicates that with our value expansion, an additional 8.5% of the hard-to-cover targets can be addressed; these targets are more challenging than the 48.5% targets covered without value expansion, as they need more contextual information to build helpful mutation hints.
Answer to RQ2: Our value expansion approach leads to a 1.5% absolute improvement in line coverage and enables us to cover 8.5% more challenging targets that require contextual mutation hints.
4.4. RQ3: Runtime Efficiency
To evaluate the runtime overhead introduced by LLM-assisted mutation, we use Number of Mutation Times and Average Execution Time per Test as metrics. Since the duration of each test is fixed at 1 hour, this metric can measure the number of mutations performed within the same time under different experimental settings. By comparing the number of mutations, we can assess the runtime overhead: the more mutations performed within the same time, the lower the time cost per mutation, indicating a smaller runtime overhead. We also calculate Average Execution Time per Test to directly measure the time cost of every run of test case mutation and evaluation. Table 4 shows the results in detail for the runtime overhead between three different configurations: the baseline, our proposed method, and our method without variable expansion (no VE).
It shows that our method causes slightly more runtime overhead than the baseline. Our method got an 18% reduction of mutation times from the baseline. In terms of average execution time per test, our method generally results in longer execution times compared to the baseline, which aligns with the intuition that the query of a language model (LLM) is expected to incur additional runtime overhead. However, the increase in runtime overhead is relatively modest, amounting to only 31% of the baseline.
In summary, our method introduces additional runtime overhead. However, this trade-off is justified by the considerable improvement in accuracy. As a result, although the number of mutations is reduced, the higher accuracy of each mutation leads to an overall increase in line coverage.